OpenStudy (anonymous):

The formula for the square area of a sphere is A=4*pi*r^2. A square sticker of side x is placed on a ball of radius 3x. What is the surface area of the sphere not covered by the sticker? write your answer in factored form.

OpenStudy (anonymous):

A=4*pi*r^2 (sphere) =4*pi*(3x)^2 =4*pi*9x^2 =pi*36x^2 or 113.01x^2 . A=x^2 (sticker) . pi*36x^2-x^2 or 112.01x^2 Surface area of the sphere not covered by the sticker. .

OpenStudy (anonymous):

thank you :)